# Feasibility-Aware Decision-Focused Learning for Predicting Parameters in the Constraints

### v1

@@ 5 Experimental Evaluation
The first is the 0-1 multi-dimensional knapsack problem (MDKP), introduced in Section 3. In the experimental section, we consider MDKP instances with 50 items and constraints in three dimensions. We consider two settings with the MDKP problem instances: 1) predicting the item weight vectors while keeping the capacity constraints known, and 2) predicting the capacity vectors with known item weights.
@@
Our third experimental setup involves the Brass alloy production problem. We reproduce this experimental setup from the work by Hu et al. [13]. It is actually a covering LP problem. We consider the data for Brass alloy, as this is publicly available in their repository  [^1] https://github.com/Elizabethxyhu/NeurIPS_Two_Stage_Predict-Optimize/. The production of the Brass alloy requires two metals – Copper (Cu) and Zinc (Zn). A factory needs to purchase ores that contain these two metals. There are 10 potential suppliers, each offering ores at a different price per unit. The exact metal content in each supplier’s ore is unknown and must be predicted. The goal is to buy a combination of ores from the 10 suppliers at the lowest possible cost, while ensuring that the total purchase contains at least 627.54 units of Cu and 369.72 units of Zn.
@@ 5.1 Dataset Description
We adapt the synthetic data generation process from the method described by Elmachtoub and Grigas [10]. This data generation approach introduces a non-linear relationship between input features and the unknown parameters. A linear model is used for training to predict the parameter using features. The motivation behind using a linear model is to demonstrate that DFL methods can yield high-quality solutions despite model misspecification. This approach has been widely used as a benchmark in several DFL studies [26, 18] for predicting parameters in the objective function. To the best of our knowledge, we are the first to extend this data generation method to the setting of predicting constraint parameters.
Each knapsack instance contains 50 items, with all items correlated with 10 input features. Item values, which appear in the objective function, are sampled independently from a Gumbel distribution with location 100 and scale 20. For each run, 1500 instances are generated, split into 900 for training, 100 for validation, and 500 for testing. We took extra caution to ensure that the synthetically-generated parameters yield non-trivial optimal solutions, avoiding cases where all or none of the items are selected in the knapsack. To achieve this, we clip the capacity in each dimension to be less than half the total weight of all items and ensure that individual item weights remain below the respective capacity values.
@@
The experiments were executed on an Intel i7-13800H (20 cores) CPU with 32GB RAM.
@@ 5.2 Experimental Results
We denote the model as Odece($\alpha$) (e.g., Odece(0.1)) to indicate Odece is trained with that specific value of $\alpha$. We compare the proposed Odece against three competitors: i) MSE: a PFL approach which trains the ML model to minimize the mse loss of the parameter predictions, ii) Comboptnet: the technique proposed by Paulus et al. [22] to compute the gradient of ILP parameters from the gradient of the loss, iii) SFL: the solver-free learning proposed by Nandwani et al. [20] for learning parameters of an ILP. For Comboptnet we compute the L1 loss of the predicted solution and the true solution and backpropagate the L1 loss through Comboptnet. We evaluate on the basis of two metrics: infeasibility and regret on the test data. For regret computation, we consider only the predicted solutions that are feasible according to the true parameters. We report normalized regret, which can be formally defined as follows:
@@
Here, $K^{\prime}$ denotes the number of feasible solutions, which varies across different techniques. It is important to note that regret can be misleading in extreme cases where most predicted solutions are infeasible, and regret is computed over only a small subset of instances.
Figure 1: Comparing infeasibility and regret of feasible solutions on Test instances (best viewed in colors).
Each technique is run five times with different seeds. We plot in Figure 1 the average proportion of infeasible solutions (x-axis) and average normalized regret (y-axis), computed over these five runs. Detailed results, including averages and standard deviations, are provided in Appendix 6. CombOptNet and SFL perform poorly on predicting weights of the MDKP problem. We looked into the gradients of these two methods and found these methods rely more on adjusting the capacity parameter of MDKP to change the predicted solution close to the true one. As a result, the gradients of the weight parameters turns out to be near zero. CombOptNet does better on predicting capacity of the MDKP and on the Brass alloy problems but still underperforms compared to Odece.
Our closest competitor is MSE, but minimizing MSE does not let decision-makers adjust the trade-off between optimality and feasibility. MSE has infeasibility rates of 33%, 60%, and 53% across the three problems respectively. For a decision-maker these infeasibility rates may be unacceptably high. Odece allows a decision-maker to attain lower infeasibility rate by varying $\alpha$, at the cost of higher normalized regret. For instance, Odece(0.8) brings the infeasibility rates down to 5%, 6%, and 14%, across the three problems respectively.
We also compared with the method by Hu et al. [13], but it exhibits high regret and infeasibility when only looking at the first stage, since it relies on a second stage to correct infeasibility and reduce regret. Because their two-stage approach differs significantly in motivation and setup from our single-stage evaluation, we omit these results from the main text. Interested readers can refer to Appendix 6 to see the performance of their approach.
@@ 6 Conclusion
We proposed Odece, a novel DFL technique for predicting parameters appearing in the constraints of COPs. Unlike existing DFL methods – which either depend on problem-specific second-stage corrective actions or which are designed to predict the full set of constraints – Odece directly optimizes for feasibility and solution quality in a single stage. Our experiments showed that by tuning Odece parameter $\alpha$, decision-makers can flexibly manage the trade-off between infeasibility and suboptimality, allowing them to align solutions with their individual preferences.
One limitation of the current Odece implementation is that it solves the COP for each instance using the predicted parameters during training. To speed up training, future work could explore caching candidate assignments, as done by Mulamba et al. [19] for predicting objective parameters. Odece could also be extended to jointly predict both constraint and objective parameters. Finally, we plan to apply Odece to a wider range of optimization problems, (such as quadratic and convex programs), as well as larger, real-world applications.
@@
This research received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation Programme (Grant No. 101002802, CHAT-Opt), from the European Union’s Horizon 2020 research and innovation programme under grant agreement No 101070149, project Tuples. and from the Research Foundation Flanders (FWO) project G0G3220N. Senne Berden is a fellow of the Research Foundation-Flanders (FWO-Vlaanderen, 11PQ024N).
@@ 18. Answer: [Yes]
  19. Justification: We are proving the code and data to reproduce the main experimental results of the paper.
@@ 33. Answer: [Yes]
  34. Justification: In the main text, we report only average of five runs for all the techniques. But we report standard deviation and conduct statistical significance in Appendix 6.
@@ 80. Guidelines:
Appendix A Problem and Dataset Description
A.1 Multi-dimensional knapsack problem
The formulation of the optimization problem is as follows:
A.1.1 Data Generation for Unknown Weight Vectors
We generated the capacity $\rho_{i}^{(\kappa)}$, the the following ways:
Infeasibility Regret
Avg Sd Avg Sd
Model
MSE 0.333 0.016 0.063 0.003
CombOptNet 0.996 0.003 0.000 0.000
SFL 0.978 0.038 0.025 0.055
TwoStageIntOpt 0.995 0.003 0.000 0.000
Odece(0.2) 0.949 0.023 0.017 0.002
Odece(0.3) 0.842 0.024 0.024 0.001
Odece(0.4) 0.595 0.064 0.039 0.003
Odece(0.5) 0.130 0.029 0.081 0.006
Odece(0.6) 0.067 0.013 0.116 0.006
Odece(0.7) 0.066 0.027 0.125 0.014
Odece(0.8) 0.045 0.009 0.139 0.007
A.1.2 Data Generation for Unknown Capacity Vectors
A.2 Brass alloy production problem
Infeasibility Regret
Avg Sd Avg Sd
Model
MSE 0.604 0.208 0.207 0.048
CombOptNet 0.721 0.046 0.289 0.053
SFL 0.442 0.244 0.374 0.129
Odece(0.2) 0.818 0.071 0.159 0.028
Odece(0.3) 0.481 0.276 0.245 0.049
Odece(0.4) 0.100 0.069 0.456 0.047
Odece(0.5) 0.070 0.030 0.474 0.114
Odece(0.6) 0.072 0.063 0.537 0.066
Odece(0.7) 0.063 0.032 0.530 0.112
Odece(0.8) 0.059 0.036 0.562 0.125
Infeasibility Regret
Avg Sd Avg Sd
Model
MSE 0.529 0.008 0.169 0.004
SFL 0.979 0.005 0.022 0.005
CombOptNet 0.321 0.136 0.293 0.101
TwoStageIntOpt (p=0.25) 0.600 0.548 2.060 0.609
TwoStageIntOpt (p=0.5) 0.758 0.433 0.762 0.985
Odece(0.2) 0.811 0.039 0.054 0.014
Odece(0.3) 0.716 0.021 0.112 0.013
Odece(0.4) 0.594 0.124 0.204 0.058
Odece(0.5) 0.494 0.065 0.244 0.037
Odece(0.6) 0.364 0.063 0.324 0.039
Odece(0.7) 0.225 0.046 0.444 0.064
Odece(0.8) 0.141 0.035 0.511 0.062
Appendix B Additional Experimental Results
For the MDKP weight prediction (Table 1) and Brass alloy production problems (Table 3), we also evaluated the TwoStageIntOpt approach [13]. However, we excluded it from the main text because it performs poorly when evaluated using single-stage metrics such as infeasibility and regret. This might be due to the fact that their method is designed for settings involving a second-stage decision problem, making its application in a single-stage context potentially unfair. Additionally, their approach targets prediction of parameters on the left-hand side of an (MI)ILP, e.g., item weights in a knapsack problem. Their implementations is not applicable when right-hand side parameters, such as knapsack capacity, are predicted. For the Brass alloy production problem, we experimented with two different penalty values included in the dataset made available by the authors. Their approach exhibited unstable learning behavior, manifested in high standard deviations across the five runs. For the knapsack weight prediction task, we fixed the penalty at 1 for all instances to maintain consistency in evaluation. Their approach exhibits more than 99% infeasibility ratio. This does not change even after changing the penalty values.
Test of statistical significance.
$\alpha$ p-value (infeasibility) p-value (regret)
0.8 $2.97\times 10^{-6}$ 0.999
0.7 $3.44\times 10^{-5}$ 0.999
0.6 $5.25\times 10^{-6}$ 0.999
0.5 $4.59\times 10^{-5}$ 0.999
0.4 0.999 $8.62\times 10^{-5}$
0.3 0.999 $4.19\times 10^{-6}$
0.2 0.999 $2.88\times 10^{-6}$
$\alpha$ p-value (infeasibility) p-value (regret)
0.8 0.003 0.995
0.7 0..03 0.995
0.6 0.006 0.999
0.5 0.002 0.997
0.4 0.002 0.999
0.3 0.228 0.855
0.2 0.982 0.040
$\alpha$ p-value (infeasibility) p-value (regret)
0.8 $4.48\times 10^{-6}$ 0.999
0.7 $6.80\times 10^{-6}$ 0.999
0.6 0.002 0.999
0.5 0.137 0.993
0.4 0.848 0.882
0.3 0.999 $2.6\times 10^{-4}$
0.2 0.999 $6.6\times 10^{-6}$
Figure 2: Validation Infeasibility vs. Epoch for MDKP Weight Prediction.
Figure 3: Validation Regret vs. Epoch for MDKP Weight Prediction.
Figure 4: Validation Infeasibility vs. Epoch for MDKP Capacity Prediction.
Figure 5: Validation Regret vs. Epoch for MDKP Capacity Prediction.
Figure 6: Validation Infeasibility vs. Epoch on the Alloy Production Problem.
Figure 7: Validation Regret vs. Epoch on the Alloy Production Problem.
Learning curves.
We also include learning curves to show how infeasibility and regret evolve over training epochs. For each problem, we present separate plots for the proportion of infeasible instances and the normalized regret, making it easier to observe the trade-off dynamics during training. Figures 2 and 3 display these trends for the MDKP problem when predicting item weights. Similarly, Figures 4 and 5 show the results for MDKP when predicting knapsack capacities. Finally, Figures 6 and 7 present the learning curves for the brass alloy production problem. These plots collectively illustrate how different methods manage the balance between feasibility and optimality throughout the training process.
Appendix C Details about Hyperparameter Configuration
Finally, in Table 7, we provide the hyperparameter configurations used for each model to reproduce the results reported above. These configurations are also reported in the Supplementary file, along with instructions for running the experiments.
Table 7: Hyperparameter configurations for each model and task.
Task Hyperparameter MSE Odece CombOptNet SFL
MDKP Weight Prediction Learning Rate 0.05 0.05 0.05 0.05
$\tau$ (CombOptNet) 0.50
Temperature (SFL) 0.50
MDKP Capacity Prediction Learning Rate 0.01 0.01 0.01 0.01
$\tau$ (CombOptNet) 0.1
Temperature (SFL) 0.1
Brass Alloy Learning Rate 0.001 0.001 0.005 0.001
$\tau$ (CombOptNet) 0.5
Temperature (SFL) 0.5

### v2

@@ 5 Experimental Evaluation
The first is the 0-1 multi-dimensional knapsack problem (MDKP), introduced in Section 3. In the experimental section, we consider MDKP instances with 50 items and constraints in three dimensions. We consider two settings: 1) predicting the item weight vectors while keeping the capacity constraints known, and 2) predicting the capacity vectors with known item weights.
@@
Our third experimental setting involves the Brass alloy production problem. We reproduce this experimental setup from the work by Hu et al. [15]. It is a covering LP problem. We use the publicly available Brass alloy data from their repository. [^1] https://github.com/Elizabethxyhu/NeurIPS_Two_Stage_Predict-Optimize/ The production of the Brass alloy requires two metals: Copper (Cu) and Zinc (Zn). A factory needs to purchase ores that contain these two metals. There are 10 potential suppliers, each offering ores at a different price per unit. The exact metal content in each supplier’s ore is unknown and must be predicted. The goal is to buy a combination of ores from the 10 suppliers at the lowest possible cost, while ensuring that the total purchase contains at least 627.54 units of Cu and 369.72 units of Zn.
@@ 5.1 Dataset Description
We adopt the synthetic data generation process described by Elmachtoub and Grigas [12]. This data generation approach introduces a non-linear relationship between input features and the unknown parameters. A linear model is used to predict the parameters from features. The motivation behind using a linear model is to demonstrate that DFL methods can produce high-quality solutions despite model misspecification. This approach has been widely used as a benchmark in DFL studies [12, 20, 22, 30, 31].
Each knapsack instance contains 50 items, with all items correlated with 10 input features. Item values, which appear in the objective function, are sampled independently from a Gumbel distribution with location 100 and scale 20. For each run, 1500 instances are generated, split into 900 for training, 100 for validation, and 500 for testing. We took extra caution to ensure that the synthetically-generated parameters produce non-trivial optimal solutions, avoiding cases where all or none of the items are selected in the knapsack. To achieve this, we clip the capacity in each dimension to be less than half the total weight of all items and ensure that individual item weights remain below the respective capacity values.
@@
The experiments were executed on an Intel i7-13800H (20 cores) CPU with 32GB RAM. Details on the problems and datasets used in the three experiments are provided in Appendix A. The source code for reproducing our results is publicly available at: https://github.com/JayMan91/OdeceDFLforConstraintsNeurips25.
@@ 5.2 Experimental Results
We denote the model as Odece($\alpha$) (e.g., Odece(0.1)) to indicate Odece is trained with that specific value of $\alpha$. We compare the proposed Odece against four competitors: i) MSE: a PFL approach which trains the ML model to minimize the MSE loss over the parameter predictions, ii) CombOptNet: the technique proposed by Paulus et al. [26] to compute the gradient of ILP parameters, iii) SFL: the solver-free learning proposed by Nandwani et al. [24] for learning parameters of an ILP, iv) 2sPtO: the two-stage predict+optimize approach proposed by Hu et al. [15]. We implement 2sPtO using a high penalty factor, as recommended in Appendix A.2 of their paper. We could not apply 2sPtO to predict the knapsack capacity, as it is designed to predict only the left-hand side parameters in linear constraints. More specifically, it can compute gradients with respect to $b$ in constraints of the form $b^{\top}\bm{x}\leq c$, but not with respect to $c$. For CombOptNet we compute the L1 loss of the predicted solution and the true solution and backpropagate the L1 loss through CombOptNet.
We evaluate performance using two metrics: the proportion of infeasible solutions and the normalized regret on the test data. Each technique is run five times with different random seeds. Figure 1 reports the average proportion of infeasible solutions (x-axis) and average normalized regret (y-axis) across these runs. Detailed results, including averages and standard deviations across the five runs, are provided in Appendix B. We compute regret only over predicted solutions that are feasible under the true parameters. More formally, out of $K$ test instances, let $K^{\prime}$ denote the number of instances where the solution $\bm{x}^{\star}\left(\bm{q}_{\kappa},\bm{\hat{\rho}}_{\kappa}\right)$ is feasible under $\bm{\rho}_{\kappa}$. The proportion of infeasible solutions is then $\tfrac{K^{\prime}}{K}$. The normalized regret is defined as follows:
@@
Note that normalized regret can be misleading in this setting. This is because in extreme cases, where most predicted solutions are infeasible, regret is computed over only a small subset of instances.
Figure 1: Infeasibility ratio and regret of feasible solutions on Test instances (best viewed in colors).
Results.
The first observation from Figure 1 is that CombOptNet exhibits relatively poor performance across the three problems. For both MDKP weight prediction and the Brass Alloy problem, the proportion of infeasible solutions generated by CombOptNet is very high ($85\%$). In MDKP capacity prediction, the proportion of infeasible solutions is slightly lower but still high ($65\%$). The performance of SFL is similar to CombOptNet for MDKP weight prediction and the Brass Alloy problem. However, in the MDKP capacity prediction problem, SFL performs significantly better, with the proportion of infeasible solutions around $25\%$. We suspect this is because SFL relies more on $c$ than $b$ to separate the true optimal solution from the negative assignments, resulting in more informative gradients for $c$ and less informative gradients for $b$. (For example, in a 3-dimensional MDKP where the optimal solution is $[0,0,1]$, negative samples such as [1,1,1][1,1,1],$[1,1,0]$, $[0,1,1]$ and $[1,0,1]$ can be classified as negative based on the capacity, i.e., the right-hand side parameter, $c$.)
For MDKP weight prediction, 2sPtO performs very poorly (the proportion of infeasible solutions is greater than $99\%$). It is not applicable for MDKP capacity prediction. Figure 1 might suggest that 2sPtO achieves a relatively lower proportion of infeasible solutions for the Brass Alloy problem, but this is misleading for two reasons. First, it results in extremely high normalized regret ($850\%$). Note that we have to break the y-axis in Figure 1 to accommodate 2sPtO’s regret values. More importantly, its performance is highly unstable, which is not noticeable in Figure 1, where the average of five runs is plotted. Table 3 in the Appendix reveals that the standard deviation across the five runs is very high. In some runs, the proportion of infeasible solutions is near 0, while in others it reaches $\approx 99\%$. This indicates unstable learning.
Our closest competitor is MSE, but minimizing MSE does not let decision-makers adjust the trade-off between optimality and feasibility. MSE has infeasibility rates of 33%, 60%, and 53% across the three problems, respectively. For a decision-maker, these infeasibility rates may be unacceptably high. Odece allows a decision-maker to attain lower infeasibility rate by varying $\alpha$, at the cost of higher normalized regret. For instance, Odece(0.8) brings the infeasibility rates down to 6%, 4%, and 18%, across the three problems, respectively. We have also conducted statistical significance tests to determine whether the test regret and infeasibility of Odece are lower than those of MSE. Since each run was conducted using the same random seed for both models, we used a paired t-test with the alternative hypothesis that MSE has higher regret and higher infeasibility than Odece. The p-values from the paired t-tests for different values of $\alpha$ are reported in Table 4, 5 and 6 in Appendix B. The results of the statistical significance test reveal that for $\alpha>0.5$, Odece produces a statistically significantly lower proportion of infeasible solutions than MSE. On the other hand, for $\alpha<0.3$, the normalized regret on feasible instances generated by Odece is statistically significantly lower than that of MSE. This holds true across all three problems.
@@ 6 Conclusion
We proposed Odece, a novel DFL technique for predicting parameters appearing in the constraints of COPs. Unlike existing DFL methods for constraint predictions – which either depend on problem-specific second-stage corrective actions or which are designed to predict the full set of constraints – Odece directly optimizes for feasibility and solution quality in a single stage. Our experiments showed that by tuning Odece’s $\alpha$ parameter, decision-makers can flexibly manage the trade-off between infeasibility and suboptimality, allowing them to align solutions with their individual preferences.
One limitation of the current Odece implementation is that it solves the COP for each instance using the predicted parameters during training. To speed up training, future work could explore caching candidate assignments, as done by Mulamba et al. [23] for predicting objective parameters. Odece could also be extended to jointly predict both constraint and objective parameters. Finally, we plan to apply Odece to a wider range of optimization problems, such as combinatorial problems with non-linear objective and constraint functions, as well as larger, real-world applications.
@@
This research received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 Research and Innovation Programme (Grant No. 101002802, CHAT-Opt), and from the Research Foundation Flanders (FWO) project G0G3220N. Senne Berden is a fellow of the Research Foundation-Flanders (FWO-Vlaanderen, 11PQ024N).
Appendix A Problem and Dataset Description
A.1 Multi-dimensional knapsack problem
The formulation of the optimization problem is as follows:
A.1.1 Data Generation for Unknown Weight Vectors
We generated the capacity $\rho_{i}^{(\kappa)}$, the the following ways:
Infeasibility Regret
Avg Sd Avg Sd
Model
MSE 0.335 0.015 0.063 0.003
CombOptNet 0.996 0.003 0.000 0.000
SFL 0.992 0.004 0.010 0.013
2sPtO 0.996 0.003 0.000 0.000
Odece(0.2) 0.952 0.020 0.017 0.002
Odece(0.3) 0.845 0.019 0.024 0.002
Odece(0.4) 0.570 0.050 0.039 0.002
Odece(0.5) 0.134 0.023 0.081 0.004
Odece(0.6) 0.079 0.027 0.110 0.009
Odece(0.7) 0.066 0.021 0.126 0.011
Odece(0.8) 0.065 0.025 0.127 0.007
A.1.2 Data Generation for Unknown Capacity Vectors
A.2 Brass alloy production problem
Infeasibility Regret
Avg Sd Avg Sd
Model
MSE 0.588 0.024 0.171 0.010
CombOptNet 0.647 0.092 0.202 0.024
SFL 0.251 0.124 0.269 0.044
Odece(0.2) 0.791 0.077 0.123 0.022
Odece(0.3) 0.572 0.269 0.175 0.086
Odece(0.4) 0.153 0.184 0.330 0.097
Odece(0.5) 0.053 0.020 0.392 0.039
Odece(0.6) 0.049 0.024 0.412 0.058
Odece(0.7) 0.055 0.020 0.417 0.042
Odece(0.8) 0.038 0.008 0.428 0.042
Infeasibility Regret
Avg Sd Avg Sd
Model
MSE 0.532 0.005 0.169 0.004
CombOptNet 0.896 0.233 0.181 0.002
SFL 0.991 0.013 0.019 0.001
2sPtO 0.400 0.548 8.539 6.691
Odece(0.2) 0.790 0.064 0.081 0.033
Odece(0.3) 0.729 0.078 0.123 0.025
Odece(0.4) 0.571 0.085 0.195 0.041
Odece(0.5) 0.471 0.123 0.270 0.054
Odece(0.6) 0.324 0.042 0.365 0.045
Odece(0.7) 0.276 0.026 0.406 0.037
Odece(0.8) 0.174 0.029 0.486 0.037
Appendix B Additional Experimental Results
Test of statistical significance.
$\alpha$ p-value (infeasibility) p-value (regret)
0.8 $3\times 10^{-6}$ 0.999
0.7 $3\times 10^{-5}$ 0.999
0.6 $5\times 10^{-6}$ 0.999
0.5 $5\times 10^{-5}$ 0.999
0.4 0.999 $1.50\times 10^{-4}$
0.3 0.999 $4\times 10^{-6}$
0.2 0.999 $3\times 10^{-6}$
$\alpha$ p-value (infeasibility) p-value (regret)
0.8 $6\times 10^{-7}$ 0.995
0.7 $2\times 10^{-7}$ 0.995
0.6 $9\times 10^{-7}$ 0.999
0.5 $1\times 10^{-6}$ 0.997
0.4 0.003 0.999
0.3 0.450 0.545
0.2 0.997 0.040
$\alpha$ p-value (infeasibility) p-value (regret)
0.8 $4\times 10^{-6}$ 0.999
0.7 $7\times 10^{-6}$ 0.999
0.6 0.002 0.999
0.5 0.160 0.993
0.4 0.848 0.882
0.3 0.999 $0.007$
0.2 0.999 $0.002$
Appendix C Details about Hyperparameter Configuration
Finally, in Table 7, we provide the hyperparameter configurations used for each model to reproduce the results reported above.
Table 7: Hyperparameter configurations for each model and task.
Task Hyperparameter MSE Odece CombOptNet SFL 2sPtO
MDKP Weight Prediction Learning Rate 0.05 0.05 0.05 0.05 0.05
$\tau$ (CombOptNet) 0.50
Temperature (SFL) 0.50
damping (2sPtO) 0.01
thr (2sPtO) 0.1
MDKP Capacity Prediction Learning Rate 0.005 0.005 0.005 0.01
$\tau$ (CombOptNet) 0.1
Temperature (SFL) 0.1
Brass Alloy Learning Rate 0.001 0.001 0.005 0.001 0.05
$\tau$ (CombOptNet) 0.5
Temperature (SFL) 0.5
damping (2sPtO) 0.01
thr (2sPtO) 0.1
@@ 18. Answer: [Yes]
  19. Justification: We are providing the code and data to reproduce the main experimental results of the paper.
@@ 33. Answer: [Yes]
  34. Justification: In the main text, we report only average of five runs for all the techniques. But we report standard deviation and conduct statistical significance in Appendix B.